The region bounded by the curve y=x²-2 and the line y=-2 and x=2 is rotated about the line y=-2. Calculate the volume of the solid formed.


1
Expert's answer
2021-04-26T02:03:14-0400

graph: y= x2-2, y=-2, x=2

r= x2-2-(-2): r=x2-4

v= "\prod\int"ba r2dx

= "\prod\int"(x2-4)2dx

=16"\prod"/9